# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v2

@@ VI Conclusion and outlook
We have proposed a generic spin-$1/2$ quasiperiodic (QP) system that unifies the existing important 1D QP models, some of which requiring an extra Majorana representation, and established a unified framework for all fundamental localization phases in quasiperiodic systems. The framework is built on three universal results. First, we showed with renormalization group method the criteria for obtaining the pure phases, namely when the system has no mobility edges. Second, we uncovered a new universal mechanism for the emergence of critical states, as characterized by the generalized incommensurate zeros in matrix elements that remain invariant under the dual transformation and introduce critical orbitals into the system. Third, we identified the condition for the exact solvability of QP systems from local constraint, which also provides the mechanism to realize effective unbounded on-site potential in spinful QP chains. These results establish a unified theory which enables an in-depth exploration of the broad classes of all fundamental localization phases in quasiperiodic systems, and offer the key insights for constructing their exactly solvable models. In particular, we proposed a novel model which hosts all basic types of mobility edges (MEs), and further predict a highly nontrivial quasiperiodic system which hosts all the seven fundamental phases in Anderson localization physics. Finally, we proposed experimentally feasible schemes to realize the theoretical predictions, both in Rydberg atom systems and in optical Raman lattice.
This work enables a broad exploration and exact characterization of the fundamental quantum phases in Anderson localization and opens up intriguing topics for the future study. For example, while the present study is focused on the spin-$1/2$ QP systems, it is natural to extend the current study to the QP systems with larger spins, e.g. to the 1D $\mathrm{SU}(N)$ QP chain with $N$ larger than $2$. Generalizing the mechanism of IDZs in matrix elements for critical states to the $\mathrm{SU}(N)$ QP systems is of great interests to obtain the rigorous critical states in generic QP systems. The exact characterization of the distinct localization phases in such $\mathrm{SU}(N)$ QP system could potentially be connected to the exactly solvable mosaic lattice models (Wang et al., 2020a; Zhou et al., 2023; Longhi, 2024a) with the large unit cell, and deserves particular future efforts. On the other hand, extending the present results to higher dimensions is another important issue for the future study.
@@ Acknowledgements.
Appendix A Proof of criteria for the pure phases without MEs
@@
The system exhibits pure phases without MEs correspond to the case that, the coefficients associated with the relevant dispersions are energy independent. In the presence of the commensurate approximation, then the band dispersion is periodic with respect to both $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. Here, $k_{x}$ is the twisted momentum attached to the hopping coupling matrix $\Pi_{j}\rightarrow\Pi_{j}e^{ik_{x}}$ and $k_{y}$ is the phase offset in the QP modulation $V_{j}^{\mathrm{d}}$ and $V_{j}^{\mathrm{od}}$. Since $P(E)$ has only real roots as $H$ is Hermitian, so the dispersion and $P(E;\kappa_{x},\kappa_{y})$ shares similar periodicity, thus $P(E;\varphi,\kappa)$ is also periodic with respect to $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$.
As indicated in the main text, we first diagonalize the on-site matrix $M_{j}$ as $M_{j}=m_{j}^{z}\sigma_{z}+m_{j}^{0}\sigma_{0}$, and the resulted hopping coupling matrix is still in the form $\Pi_{j}=\sum_{s}p_{j}^{s}\sigma_{s}$, with $s$ being $s=\{0,x,y,z\}$. For the criteria that the coefficients of hopping coupling matrix satisfy
@@
and both coupling matrices exclude $\sigma_{0}$, namely
$\Pi_{j}=p_{j}^{x}\sigma_{x}+p_{j}^{y}\sigma_{y}+p_{j}^{z}\sigma_{z},\quad M_{j}=m_{j}^{z}\sigma_{z}.$ (60)
We can first simplify model by denoting $m_{j}^{z}\equiv m_{j}$ and
$p_{j}^{x}=p_{j}^{\bot}\cos\theta,\quad p_{j}^{x}=p_{j}^{\bot}\sin\theta,$ (61)
@@
then the system becomes
@@
and $\vec{v}_{j}=U(z,\theta)\vec{u}_{j}=\left(\begin{array}[]{cc}v_{j\uparrow}&v_{j\downarrow}\end{array}\right)^{\intercal}$.
Before proving the absence of the MEs when satisfying the criteria, we first identify the chiral symmetry of the system under this condition, which is highly related to the disappearance of the MEs. We notice that under the unitary operator ${\cal O}=\sigma_{y}$, both coupling matrices change sign, i.e.,
${\cal O}\Pi^{\prime}_{j}{\cal O}^{\dagger}=-\Pi^{\prime}_{j},\quad{\cal O}M_{j}{\cal O}^{\dagger}=-M_{j}.$ (65)
Therefore both $\vec{v}_{j}$ and ${\cal O}\vec{v}_{j}$ are the eigenstate of the system, with the corresponding energies come in pair as $(E,-E)$.
@@
By the commensurate approximation, we approximate the irrational number $\alpha$ by the commensurate size, then $L$ becomes the size of the unit cell and the system exhibits the periodic structure. Then to calculate the characteristic polynomial $P(E)=\det|H-EI|$, we can perform the Fourier transform, i.e. we multiply by $e^{i2\pi\alpha jm}$ and sum over $j$, then finding characteristic polynomial $P(E)=\det|H-EI|$ becomes finding the determinant $P(E)=\det|A(E)|$, with $A(E\text{)}$ being
@@
This form has the advantages that within each element, the energy term $E$ is decoupled from the hopping $p_{j}$ and on-site term $m_{j}$, facilitating the discussion.
To prove that in this case, the system exhibits pure phase with MEs, we are going to prove that the effective coefficients of the dispersion as the we iterating the commensurate approximated system size $L$ will saturate to a energy-independent parameter. First of all, due to the presence of the chiral symmetry, the spectrum of the system is symmetric in energy $E$ with respective to zero energy. Therefore the characteristic polynomial $P(E)=\det|A|$ can only be the polynomial of energy $E$ with even power. In addition, due to the periodic structure of the Hamiltonian, the characteristic polynomial is the function of $Lk_{x}$ and $Lk_{y}$. Now we consider the case of odd $L$.
@@
Appendix B Exact solvability of QP systems from local constraint

### v4

@@ VI Conclusion and outlook
We have proposed a generic spin-$1/2$ quasiperiodic (QP) system that unifies the existing important 1D QP models, some of which requiring an extra Majorana representation, and established a unified framework for all fundamental localization phases in quasiperiodic systems. The framework is built on three universal results. First, we showed with renormalization group method the criteria for obtaining the pure phases, namely when the system has no mobility edges. Second, we uncovered a new universal mechanism for the emergence of critical states, as characterized by the generalized incommensurate zeros in matrix elements that remain invariant under dual transformation and introduce critical orbitals into the system. Third, we identified the condition for the exact solvability of QP systems from local constraint, which also provides the mechanism to realize effective unbounded on-site potential in spinful QP chains. These results establish a profound unified theory for all fundamental localization phases in quasiperiodic systems, and offer the key insights for constructing their exactly solvable models. In particular, we proposed the novel spin-selective QP lattice model and QP optical Raman lattice model, in which we predict all basic types of mobility edges (MEs) and all the seven fundamental phases in Anderson localization physics, respectively, which have not been achieved anywhere else. Finally, the experimentally feasible schemes to realize the theoretical predictions have been proposed and studied in detail.
This work enables a broad exploration and exact characterization of the fundamental quantum phases in Anderson localization and opens up intriguing topics for the future study. For example, while the present study is focused on the spin-$1/2$ QP systems, it is natural to extend the current study to the QP systems with larger spins, e.g. to the 1D $\mathrm{SU}(N)$ QP chain with $N$ larger than $2$. Generalizing the mechanism of IDZs in matrix elements for critical states to the $\mathrm{SU}(N)$ systems is of great interests to obtain the rigorous critical states in generic QP systems. The exact characterization of the distinct quantum phases in such $\mathrm{SU}(N)$ QP system could potentially be connected to the exactly solvable QP mosaic models (Wang et al., 2020a; Zhou et al., 2023) with large unit cells, and deserves particular future efforts. On the other hand, extending the present results to higher dimensions is another important issue for the future study.
@@ Acknowledgements.
Appendix A 1D spinless QP model
In this section, we show that the 1D spinless QP model with nearest-neighbor hopping can also be unified within the spinful QP framework, under the Majorana representation. For a generic 1D spinless QP chain model nearest-neighbor hopping, the tight binding Hamiltonian is given by
The $t_{j}$ and $V_{j}$ are the hopping coefficients and on-site potential, respectively. Such 1D spinless QP chain can be rewritten into the spinful QP chain by introducing the Majorana representation
$c_{j}=\frac{1}{2}(\gamma_{B,j}+i\gamma_{A,j}),\ c_{j}^{\dagger}=\frac{1}{2}(\gamma_{B,j}-i\gamma_{A,j}),$ (65)
with $\gamma_{\sigma,j}=\gamma_{\sigma,j}^{\dagger}$ and $\{\gamma_{\sigma,j},\gamma_{\sigma;,j^{\prime}}\}=2\delta_{\sigma\sigma^{\prime}}\delta_{jj^{\prime}}$. In this basis, the Hamiltonian in Eq. 64 becomes
$+$ $\sum_{j}\frac{V_{j}}{2}(1+i\gamma_{B,j}\gamma_{A,j}).$
$\Pi_{j}=\frac{t_{j}}{4}\sigma_{y},\ M_{j}=\frac{V_{j}}{4}\sigma_{y}.$ (67)
One characteristic is that both hopping coupling and on-site matrix share the same Pauli component $\sigma_{y}$, which is a hallmark of the spinless QP model rewritten in the spinful QP framework. This formalism unifies all the important 1D spinless QP model in this context, including the AA model Aubry and André (1980), the extended AA (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015), the generalized AA (GAA) model Ganeshan et al. (2015), or the GPD model in some context. These are summarized in the table. 3.
Models $t_{j}$ $V_{j}$
AA model $t$ $2V_{0}V_{j}^{\mathrm{d}}$
EAA model $(t+\mu)V_{j}^{\mathrm{od}}$ $2V_{0}V_{j}^{\mathrm{d}}$
GAA model $t$ $2V_{0}V_{j}^{\mathrm{d}}/(1-aV_{j}^{\mathrm{d}})$
Table 3: The parameters of different 1D spinless QP model
The conditions for the different states of the AA, EAA and GAA models are summarized in Table 4.
Extended Critical Localized
AA $t>V_{0}$ $t=V_{0}$ $t<V_{0}$
EAA $t>{\rm max}\{\mu,V_{0}\}$ $\mu>{\rm max}\{t,V_{0}\}$ $V_{0}>{\rm max}\{t,\mu\}$
GAA $aE<2(t-V_{0})$ $aE=2(t-V_{0})$ $aE>2(t-V_{0})$
Table 4: The criteria for different states of AA, EAA and GAA models. Here $E$ is the eigenenergies of the system. Without loss of generality, we assume of the hopping and on-site coefficients are positive.
Appendix B Proof of criteria for the pure phases without MEs
@@
The system exhibits pure phases without MEs correspond to the case that, the coefficients associated with the relevant dispersions are energy independent. In the presence of the commensurate approximation, then the band dispersion is periodic with respect to both $\kappa_{x}=Lk_{x}$ and $\kappa_{y}=Lk_{y}$. Here, $k_{x}$ is the twisted momentum attached to the hopping coupling matrix $\Pi_{j}\rightarrow\Pi_{j}e^{ik_{x}}$ and $k_{y}$ is the phase offset in the QP modulation $V_{j}^{\mathrm{d}}$ and $V_{j}^{\mathrm{od}}$. Since $P(E)$ has only real roots as $H$ is Hermitian, so the dispersion and $P(E;\kappa_{x},\kappa_{y})$ shares similar periodicity, thus $P(E;\varphi,\kappa)$ is also periodic with respect to $\kappa_{x}$ and $\kappa_{y}$.
Before proving the absence of the MEs when satisfying the criteria in the main text, we first establish the condition that enforces chiral symmetry. In the presence of chiral symmetry, one readily shows that the Hamiltonian excludes any spin-independent, i.e. $\sigma_{0}$ term. We then diagonalize the on-site matrix $M_{j}$ as $M_{j}=m_{j}^{z}\sigma_{z}$, and the resulted hopping coupling matrix retains the form $\Pi_{j}=\sum_{s}p_{j}^{s}\sigma_{s}$, with $s$ being $s=\{x,y,z\}$. When the hopping coupling matrix satisfy
@@
then the system exhibits the chiral symmetry. We simplify notation by setting $m_{j}^{z}\equiv m_{j}$. And Due to the condition in Eq. (69), the transverse hopping coupling components can be parameterized as
$p_{j}^{x}=p_{j}^{\bot}\cos\theta,\ p_{j}^{x}=p_{j}^{\bot}\sin\theta,$ (70)
@@
then we have the eigenvalue equation
@@
and $\vec{v}_{j}=U(z,\theta)\vec{u}_{j}=(v_{j\uparrow},v_{j\downarrow})^{\intercal}$. Thus, under the unitary operator ${\cal O}=\sigma_{y}$, both coupling matrices anticommute
${\cal O}\Pi^{\prime}_{j}{\cal O}^{\dagger}=-\Pi^{\prime}_{j},\ {\cal O}M_{j}{\cal O}^{\dagger}=-M_{j}.$ (74)
Therefore both $\vec{v}_{j}$ and ${\cal O}\vec{v}_{j}$ are the eigenstate of the system, with the corresponding energies come in pair as $(E,-E)$, confirming the chiral symmetry.
@@
By the commensurate approximation, we approximate the irrational number $\alpha$ by the commensurate size, then $L$ becomes the size of the unit cell and the system exhibits the periodic structure. Then to calculate the characteristic polynomial $P(E)=\det|H-EI|$, we can perform the Fourier transformation, i.e. we multiply by $e^{i2\pi\alpha jm}$ and sum over $j$, then finding characteristic polynomial $P(E)=\det|H-EI|$ becomes finding the determinant $P(E)=\det|A(E)|$, with $A(E\text{)}$ being
@@
As mentioned in the main text, this structure ensures the energy terms and momentum dispersions decouple naturally, eliminating cross terms such as $Ee^{ik_{x}}$. Our task is to separately show that the dominant dispersions for extended, localized, and critical states have energy-independent coefficients.
@@
Appendix C Generalized 2D bilayer mapping
In this section, we provide the detail of the mapping the spinful QP model to a generalized 2D bilayer square lattice system, in the presence of an external magnetic field with an irrational flux threading each unit cell. We provide the detail that the uniform and QP hopping couplings in $\Pi_{j}$ are mapped to the hopping couplings along $x$-direction and diagonal directions, respectively. And the uniform and QP on-site coupling in $M_{j}$ are mapped to on-site coupling and hopping along $y$-direction, respectively. This mapping is done by interpreting the phase shift $k_{y}$ as the quasi-momentum in the y-direction and the spin index $s$ as the layer index. Performing a Fourier transformation, the coefficient $V_{j}^{\rm od}$ in hopping coupling matrix $\Pi_{j}$ and $V_{j}^{\rm d}$ in the on-site matrix $M_{j}$ contribute the hopping terms along diagonal and y-directions,respectively. In the following, we analyze this generalized mapping of different processes.
Under the 2D mapping, the uniform hopping coupling in the Hamiltonian is mapped to the intra-layer (for $s=s^{\prime}$) or inter-layer (for $s\neq s^{\prime}$) hopping coupling in the $x$-direction
$t^{ss^{\prime}}(c_{j+1s}^{\dagger}c_{js^{\prime}}+{\rm h.c.})\rightarrow t^{ss^{\prime}}(a_{x+1,y,s}^{\dagger}a_{x,y,s^{\prime}}+{\rm h.c.}),$ (78)
with $t^{ss^{\prime}}$ being the uniform hopping coefficients of different elements in $\Pi_{j}$. The QP hopping term in the Hamiltonian is mapped to the intra-layer or inter-layer hopping coupling in the diagonal direction in the presence of magnetic flux, which is given by
$\mu^{ss^{\prime}}V_{j}^{\rm od}(c_{j+1s}^{\dagger}c_{js^{\prime}}+{\rm h.c.})$$\rightarrow$ $\frac{\mu^{ss^{\prime}}}{2}(e^{-i2\pi\alpha x}a_{x+1,y+1,s}^{\dagger}a_{x,y,s^{\prime}}$ (79)
$+$ $e^{i2\pi\alpha x}a_{x+1,y-1,s}^{\dagger}a_{x,y,s^{\prime}}$
$+$ ${\rm h.c.}),$
where $\mu^{ss^{\prime}}$ are the QP hopping coefficients. The uniform on-site coupling $\lambda^{ss^{\prime}}c_{js}^{\dagger}c_{js^{\prime}}$ is mapped to the uniform onsite term or inter-layer coupling. For $s=s^{\prime}$, the mapping yields the on-site energy shift, and when $s\neq s^{\prime}$, it gives the inter-layer coupling
$\lambda^{ss^{\prime}}c_{js}^{\dagger}c_{js^{\prime}}\rightarrow\lambda^{ss^{\prime}}(a_{x,y,s}^{\dagger}a_{x,y,s^{\prime}}+{\rm h.c.}).$ (80)
Finally the QP on-site coupling $V^{ss^{\prime}}V_{j}^{\rm d}c_{js}^{\dagger}c_{js^{\prime}}$ characterizes the intra-layer or inter-layer hopping coupling in the $y$-direction under the magnetic field, with $V^{ss^{\prime}}$ being the hopping coefficients
$V^{ss^{\prime}}V_{j}^{\rm d}c_{js}^{\dagger}c_{js^{\prime}}\rightarrow\frac{V^{ss^{\prime}}}{2}(e^{-i2\pi\alpha x}a_{x,y+1,s}^{\dagger}a_{x,y,s^{\prime}}+{\rm h.c.}).$
From the mapping given above, we can see that the dominant constant hopping term (quasiperiodic onsite term) in the original 1D model gives the dominant hopping along the $x$ ($y$) direction in the mapped 2D bilayer system, and the extended (localized) states emerge. In contrast, when the quasiperiodic hopping dominates in the 1D model, in the 2D bilayer system the inter-layer diagonal hopping coupling dominates over other terms. In this case the states are extended in the diagonal direction, hence being delocalized in both $x$ and $y$ directions.
The states delocalized in both directions are critical states, and are invariant under dual transformation. In the mapped 2D bilayer system, the dual transformation $c_{j,s}=(1/\sqrt{L})\sum_{n}e^{-i2\pi n\alpha j}c_{n,s}$ is equivalent to a gauge transformation, which is given by
$a^{\dagger}_{x,y,s}\rightarrow e^{i\varphi xy}a^{\dagger}_{x,y,s},\ \vec{A}=(-\varphi x,0)\rightarrow\vec{A}=(0,\varphi y).$ (82)
The resulted Hamiltonian after the transformation yields
$H=$ $\sum_{x,y,s,s^{\prime}}\Big{[}t^{ss^{\prime}}e^{i\varphi y}a_{x+1,y,s}^{\dagger}a_{x,y,s^{\prime}}+\frac{V^{ss^{\prime}}}{2}a_{x,y+1,s}^{\dagger}a_{x,y,s^{\prime}}$
$+\lambda^{ss^{\prime}}a_{x,y,s}^{\dagger}a_{x,y,s^{\prime}}+\frac{\mu^{ss^{\prime}}}{2}(e^{i\varphi y}a_{x+1,y+1,s}^{\dagger}a_{x,y,s^{\prime}}$
$+e^{-i\varphi y}a_{x+1,y-1,s}^{\dagger}a_{x,y,s^{\prime}})+{\rm h.c.}\Big{]}.$ (83)
It is seen that the diagonal hopping terms and on-site coupling terms in the 2D bilayer remain invariant under the dual transformation, while the hopping along the $x$ and $y$ directions are interchanged. Thus when the hopping along diagonal direction dominates, while those along $x$ and $y$ directions are irrelevant, the states are self-dual and are critical.
Appendix D Semiclassical approximation of the critical states
We take the semiclassical approximation to study how the critical states are obtained under the mechanisms of the generalized incommensurate zeros (GIZs) in hopping coupling matrix $\Pi_{j}$ and the incommensurately distributed zeros (IDZs) in the matrix component of on-site matrix $M_{j}$ shared in $\Pi_{j}$. Note that while this is an approximate study, it provides an intuitive picture for the emergence of the critical states.
Figure 7: Phase space trajectories for extended, localized and critical states. The extended state is continuous in $x$ direction while the localized state is continuous along momentum direction. The critical state is continuous in both directions. Here the extended, localized and critical states are realized in the extended AA model by setting $(t,V,\mu)=(1,2.5,0)$, $(t,V,\mu)=(1,0.5,0)$ and $(t,V,\mu)=(1,2.5,1.5)$, respectively.
The semiclassical description is well valid for small irrational number $\alpha\ll 1$, in which case the QP modulation varies slowly across the lattice. We first take the example of the extended AA model introduced in Appendix A, which supports all three types of states. Then we have the energy in the phase space
$E(p,x)=[2t+2\mu\cos(2\pi\alpha x)]\cos p+V\cos(2\pi\alpha x).$ (84)
Here $x=j$ labels the site coordinate and $p$ is the momentum, so that the contours $E(p,x)=E$ define the phase-space trajectories Albert and Leboeuf (2010); Malla and Raikh (2018) of the extended AAH model. As depicted in Fig. 7, the extended and localized states correspond to the trajectories continuous in $x$ direction but discontinuous in $p$ direction, whereas localized states exhibit the opposite behavior. The critical states are continuous in both directions. The dual transformation here is swapping $2\pi x$ and $p$. One can notice that $2\mu\cos(2\pi\alpha x)\cos p$ remains invariant under the dual transformation and always percolates in both directions, responsible for the critical states.
The semiclassical phase-space trajectory for the current spin-$1/2$ QP system decomposes into two branches,
$E_{\pm}(p,x)=h_{0}\pm\sqrt{h_{x}^{2}+h_{y}^{2}+h_{z}^{2}},$ (85)
where each component is given by
$h_{0}$ $=V_{0}\cos\bigl(2\pi\alpha x\bigr{missing})+2t_{0}\cos p+2\mu_{0}\cos\bigl(2\pi\alpha x\bigr{missing})\cos p,$ (86)
$h_{x}$ $=\lambda_{x}+V_{x}\cos\bigl(2\pi\alpha x\bigr{missing})+2|t_{x}|\cos\bigl(p+\phi_{t_{x}}\bigr{missing})$
$+2|\mu_{x}|\cos\bigl(2\pi\alpha x\bigr{missing})\cos\bigl(p+\phi_{\mu_{x}}\bigr{missing}),$ (87)
$h_{y}$ $=\lambda_{y}+V_{y}\cos\bigl(2\pi\alpha x\bigr{missing})+2|t_{y}|\cos\bigl(p+\phi_{t_{y}}\bigr{missing})$
$+2|\mu_{y}|\cos\bigl(2\pi\alpha x\bigr{missing})\cos\bigl(p+\phi_{\mu_{y}}\bigr{missing}),$ (88)
$h_{z}$ $=\lambda_{z}+V_{z}\cos\bigl(2\pi\alpha x\bigr{missing})+2t_{z}\cos p$
$+2\mu_{z}\cos\bigl(2\pi\alpha x\bigr{missing})\cos p.$ (89)
Here, $|t_{x,y}|$ and $\phi_{t_{x,y}},\phi_{\mu_{x,y}}$ denote the magnitudes and phase angles of the generally complex spin-flipped hopping coefficients, while the spin-conserved hopping coefficients are taken to be real up to a global gauge choice, and all the other parameters are real.
The critical states generated from the GIZs can be understood as the phase-space trajectory percolation due to explicit dual invariant terms. When the system exhibits GIZs in the hopping coupling matrix $\Pi_{j}$, arising from either the spin-flipped or spin-conserved QP terms, the $\cos(2\pi\alpha x)\cos p$ contribution dominates the local dispersion, yielding a percolating, dual-invariant trajectory in phase space and thus producing critical eigenstates.
The semiclassical picture also explains that the IDZs in the matrix component of $M_{j}$ shared in $\Pi_{j}$ generate critical states. Without losing generality, we consider IDZs in $\sigma_{z}$ component of $M_{j}$ shared in $\Pi_{j}$ as an example. Specifically, the uniform hopping couples to $\sigma_{z}$ component and $M_{j}$ exhibits IDZs in $\sigma_{z}$ component. This case has no GIZs for $\mu_{0,x,y,z}=0$, the $h_{z}$ takes the generic form
$h_{z}=\lambda_{z}+V_{z}\cos\bigl(2\pi\alpha x\bigr{missing})+2t_{z}\cos p,$ (90)
[…]
Appendix E Exact solvability of QP systems from local constraint